• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name J.L.Herraiz

    Log entry time 16:01:13 on December 8,2007

    Entry number 219139

    keyword=Day Shift Summary

    No beam because of Energy Change - No beam expected for the whole shift.

    The spectrometers are in the correct angle (90 degress both arms) and with the correct momentum for the next runs.

    Jack Segal asked us for Controlled Access to play with the Spectrometers (Q2 and Q3 of both arms) during this energy change.

    No beam until the end of the shift.